OpenStudy (anonymous):

I'm trying to find the square root for this (4y+6)^2=36

OpenStudy (amistre64):

divide of the 4 and subtract the 6

OpenStudy (amistre64):

or y = 0 for one

OpenStudy (amistre64):

-3 for the other?

OpenStudy (anonymous):

divide the 4 by what?

OpenStudy (amistre64):

nothing, i was reading a different problem in my head :)

OpenStudy (anonymous):

oh

OpenStudy (amistre64):

if we sqrt the sides we get: 4y+6 = +- 6

OpenStudy (anonymous):

4y+6 = 6 ===> y=0 or 4y+6 = -6 ==> y=12/4=3

OpenStudy (anonymous):

I mean y =-12/4=-3

OpenStudy (amistre64):

4y+6 = +- 6 -6 -6 ------------ 4y = 0 , -12 /4 /4 ----------- y = 0, -3